# GASLITEing the Retrieval: Exploring Vulnerabilities in Dense Embedding-based Search

### v1

@@ 6.4 The perfect Attack
Motivated to understand how well GASLITE reflects models’ susceptibility, we compute the performance of a hypothetical attack, perfect, that perfectly optimizes Eq. 2, thus simulating an optimal run of GASLITE. 333While perfect embodies the optimal vectors for Eq. 2, these are still not guaranteed to be optimal for the attack in general (§4.1), subsequently GASLITE may inadvertently converge to better solutions. To this end, we perform the aforementioned evaluations on vectors providing an optimal solutions to Eq. 2 (i.e., query centroid per cluster; see App. E.5). Note this is merely a simulation, as an actual attack needs to invert the vectors into text.
@@
First, we treat perfect’s attack success as a strong reference measure, comparing it to GASLITE’s under Knows What setting §6.2 with budget =10 (Fig. 6(a)). We observe that, while GASLITE’s success varies between models, in all cases it exhausts most of perfect’s attack success. This result shows that, under our framework, GASLITE’s performance is near-ideal.
@@
Additionally, perfect can also be used to efficiently estimate the potential attack success without running the full attack. We use this to extrapolate the attack success for prohibitively large budgets (Fig. 6(b)), observing the attack performance in §6.3 (Knows Nothing), attained with a budget of =100, can be further increased via additional adversarial passages, while maintaining a relatively low poisoning rate (e.g., 0.01% for =1K). Finally, this extrapolation further emphasizes the variability of models’ susceptibility, which we discuss next.
7 Analyzing Potential Vulnerabilities in Embedding Spaces
@@
[…]
Geometry of the Embedding Space. Each model learns an embedding space of potentially different geometry. Focusing on cosine similarity models—limiting attacks to directions within the embedding space—we observed that, for example, the E5 model was consistently more vulnerable than MiniLM (e.g., Fig. 3). Inspecting the geometry of their embedding spaces, we find E5’s, as opposed to MiniLM’s, to not be uniformly distributed (Fig. 8(a)); that is, random text pairs produce high similarities, contradicting mathematical intuition.444The expected cosine similarity of uniformly-distributed (of isotropic distribution) high-dimensional vectors is , where is their dimension (Vershynin, 2018). In the case of most evaluated models =768. This phenomenon, also observed in other evaluated models (Fig. 8(a)), is known as anisotropy of text representations (Ethayarajh, 2019). While seemingly not correlated with benign performance (Ait-Saada & Nadif, 2023),
@@
[…]
8 Limitations
This work focuses on attacking embedding-based dense retrieval. While GASLITE demonstrates improved speed and performance compared to previous attacks on text models (Fig. 2(a)) and efficient baselines (§6), it remains compute-intensive, in tandem with other discrete optimizers (more details in App. H.8). Although seemingly demanding resource-wise, increasing the attack budget () in certain scenarios (e.g., §6.3), could potentially enhance attack performance (§6.4), underscoring both the potential and risks of more efficient attacks following this work. Moreoever, like other text-domain attacks, GASLITE’s success can be mitigated by defenses such as perplexity filtering to eliminate non-fluent passages (Jain et al., 2023). Still, an effort can be made to bypass these defenses, as we demonstrate with GASLITE in App. G, and as noted in recent LLM jailbreak research (Paulus et al., 2024; Thompson & Sklar, 2024).
9 Conclusion
@@
This work highlights the potential risks in embedding-based search through our method GASLITE (§4), surpassing prior approaches (§4.2, §6). Our extensive evaluation across nine widely used retrievers and three different threat models, demonstrates embedding-based search’s susceptibility to SEO via corpus poisoning. In particular, promoting concept-specific information with GASLITE can be done efficiently (requiring 0.0001% poisoning rate) and with high success (Fig. 3, §6.2), nearing the performance of a strong hypothetical attack (perfect in §6.4). Considering other SEO settings, we find single-query attacks possible with optimal success (Tab. 1, §6.1), as opposed to indiscriminately targeting a set of diverse queries, which we find more challenging and budget-demanding, albeit possible (Fig. 5, §6.3). Furthermore, we observe some models are consistently more vulnerable to attacks than others; we identify key factors potentially affecting model susceptibility to corpus poisoning (§7), including similarity metric and the anisotropy phenomenon.
@@
Future work. Future research may explore further constraining retrieval attacks (e.g., requiring fluency), following our initial defense-bypass results (App. G). Additionally, as our formulation reduces corpus poisoning attacks to controlled embedding-inversion to text (Eq. 2), recent advances in such methods (Morris et al., 2023) may enable the development of more efficient attacks. Lastly, our insights into model susceptibility variance (§7) may provide a foundation for further exploration of embedding-based retrieval robustness.
Ethics Statement
@@
Our paper proposes a practical attack against embedding-based search via corpus poisoning, demonstrated on widely-used models. While our work aims to advance the security of embedding-based text retrieval systems, we recognize the potential for misuse. After careful consideration, we believe the benefits of publishing this research outweigh the potential risks for several reasons.

### v2

@@ 6.4. The perfect Attack
Motivated to understand how well GASLITE reflects models’ susceptibility, we compute the performance of a hypothetical attack, perfect, that perfectly optimizes Eq. 2, thus simulating an optimal run of GASLITE. 666While perfect embodies the optimal vectors for Eq. 2, these are still not guaranteed to be optimal for the attack in general (§4.1), subsequently GASLITE may inadvertently converge to better solutions. To this end, we perform the aforementioned evaluations on vectors providing an optimal solutions to Eq. 2 (i.e., query centroid per cluster; see App. E.5). Note this is merely a simulation, as an actual attack needs to invert the vectors into text.
@@
First, we treat perfect’s attack success as a strong reference measure, comparing it to GASLITE’s under Knows What setting (§6.2) with budget =10 (Fig. 6(a)). We observe that, while GASLITE’s success varies between models, in all cases it exhausts most of perfect’s attack success. This result shows that, under our framework, GASLITE’s performance is near-ideal.
@@
Additionally, perfect can also be used to efficiently estimate the potential attack success without running the full attack. We use this to extrapolate the attack success for prohibitively large budgets (Fig. 6(b)), observing the attack performance in §6.3 (Knows Nothing), attained with a budget of =100, can be further increased via additional adversarial passages, while maintaining a relatively low poisoning rate (e.g., 0.01% for =1K). Finally, this extrapolation further emphasizes the variability of models’ susceptibility, as we discuss next (§9).
6.5. Attack Generalizability
6.5.1. Generalizability to Unseen Query Sets
PPL Filtering
L2 Filtering
In the following we study cross query-set transferability under the Knows Nothing setting. Following Zhong et al. (Zhong et al., 2023) results showing potential for transferring Cor.Pois. across datasets, we experiment GASLITE under this setting. Concretely, we use GASLITE to craft adversarial passage given with source queries (e.g., a sample of MSMARCO’s train set) and evaluate poisoning with those passages another, target, dataset (e.g., FiQA2018, SciFact and Quora test sets; (Muennighoff et al., 2023)). Said differently, we evaluate the success of adversarial passages, crafted as part of the Knows All attacks (on MSMARCO and NQ; §6.3), on other datasets of different query distributions.
Fig. 7 shows that attacks generalize across query sets, consistent with the findings of Zhong et al. (Zhong et al., 2023). In particular, topic-specific query sets (FiQA and SciFact) show high susceptibility to this method. We hypothesize that the semantic diversity in the source query sets (e.g., in MSMARCO) is useful for crafting query-universal attacks that apply effectively to target query sets of narrower semantics.
6.5.2. Applicability to Additional Architectures
Deviating from our focus on popular embedding-based models that are compact and rely on encoder-only architectures (e.g., BERT (Devlin et al., 2019)), we also study embeddings produced by additional backbone architecture, including decoder-only LLMs, selecting top-ranked ranked representative models from the MTEB leaderboards (mte, [n. d.]). Specifically, we apply GASLITE and baselines against Qwen3-Embedding (0.6B; based on the LLM model Qwen3) (Zhang et al., 2025a), GTE-ModernBERT (with ModernBERT backbone) (Warner et al., 2024), and Stella (1.5B; based on the LLM model Qwen2) (Zhang et al., 2025b), with full details in App. G.4. We evaluate under the Knows All and Knows What settings, with results shown in Tab. 18 and Tab. 19; App. G.4 respectively. The results reflect trends similar to those observed for compact encoder-only models. Concretely, GASLITE consistently achieves 100% appeared@10 under Knows All for all models, similar to §6.1. Under Knows What, the susceptibility of Qwen3 and GTE-ModernBERT is comparable to the average model susceptibility evaluated in §6.2, while Stella exhibits slightly higher robustness, which we believe can be further challenged by optimizing GASLITE against models of larger scales. Overall, this confirms that GASLITE poses a threat to embedding-based retrieval models across various architectures and sizes.
7. Defenses and Adaptive Variants Against Them
We evaluate GASLITE against previously proposed detection-based baseline defenses (Jain et al., 2023; Alon and Kamfonas, 2023; Zhong et al., 2023)—specifically, perplexity-based and norm-based—and form baseline variants that adapt GASLITE for bypassing these. Future work may build upon our adaptive variants to directly improve attack stealthiness.
Concretely, in these defenses, detection aims to filter adversarial passages out of the corpus, while setting a threshold of zero false-positives (e.g., picking maximal perplexity of benign passages). To construct the adaptive variants of our attack, we add an evasion term (e.g., an LM perplexity) to the optimized adversarial objective (i.e., Eq. 2). We also consider simpler variants, involving other attack parameters. We present the technical details in App. F.4.
Setup. We evaluate the defenses and our adaptive variants under Knows What (§6.2), on two different models, picked for their high benign performance (Tab. 4 in App. E): E5 on fluency-based detection, and Contriever-MS for -norm-based detection. Notably, as we show, GASLITE’s variants attack success after applying the defenses is virtually higher than prior attacks before defenses, thus we exclude the latter from this evaluation.
Fluency-based Detection. GASLITE, similarly to many other attacks (e.g., (Zhong et al., 2023; Zou et al., 2023)), may result in non-fluent and nonsensical triggers, which can be exploited to identify adversarial passages via perplexity filtering (Jain et al., 2023; Alon and Kamfonas, 2023); indeed, such filtering can be successfully applied on GASLITE, as shown in Fig. 8(b). To bypass this defense we propose enriching the adversarial objective by adding a GPT2 perplexity term (Radford et al., 2019) (GASLITE-Flu). For simpler adaptive variants, we consider limiting the trigger length to (GASLITE10), and sampling token candidates, throughout the optimization, only from the top- logits of GPT2 (GASLITE-FluLogits). Fig. 8(a) shows that all adaptive variants preserve most attack success after applying the defense, as these assimilate in the perplexity distribution of the benign passages (Fig. 8(b)). Qualitative samples of fluent adversarial passages in Tab. 16.
-Norm-based Detection. We observe that attacking dot-product similarity models results with crafted passages of large -norm (§9); as noted by Zhong et al. 2023 (Zhong et al., 2023), this can be used to identify anomalous adversarial passages, filtering passages surpassing the maximal benign -norm. Indeed, as shown in Fig. 8(d), we find this defense to successfully filter GASLITE passages. However, enriching the adversarial objective with a term that penalizes large--norm passages (GASLITE-L2; see App. F.4 for details) significantly decreases the -norm (Fig. 8(d)), thus increasing the post-defense attack success (Fig. 8(c)). The results present a trade-off between -norm of adversarial passages and attack success, as we later discuss (§9). While we expect this trade-off can be further optimized, the variant of GASLITE-L2 we evaluate achieves substantially higher success than the baselines (reported in §E.4) after applying the -norm-based defense (9.6% vs. 0% appeared@10).
8. Case Studies
We demonstrate GASLITE’s practicality and applicability, through targeting an actual, popular RAG application (§8.1), and a search system queried for product recommendations (§8.2).
8.1. Attacking a RAG System
We target a popular chat application (OpenWebUI) that employs embedding-based search for RAG. Specifically, we investigate whether the exact adversarial passages crafted for our controlled evaluation (§6) can be effectively reused to attack a practical system. Such a system introduces new challenges, including variations in the underlying corpus and user query distributions, thus deviating from those used to train the attack. This adds to the complexities of end-to-end systems; for instance, the RAG pipeline chunks the corpus into sub-passages (i.e., divides large passages into smaller segments) (Lewis et al., 2020), a preprocessing step that often modifies the original adversarial passages used for poisoning.
Setting. We target the LLM-based, retrieval-augmented chat application, OpenWebUI (Open WebUI, [n. d.]), which allows setting-up a local chat, that accesses a customized knowledge base. We simulate a use-case of the knowledge-augmented chat by uploading the seven books of Harry Potter as the RAG textual corpus (Rowling, [n. d.]), where the user queries regard them. We use the default configuration of OpenWebUI, which includes the open-source embedding model MiniLM for retrieval, and the RAG configuration (e.g., chucking and prompt templates), and set GPT-4o-mini as the LLM (OpenAI, 2024b).777LLM was chosen arbitrarily, and the attack is similarly applicable to any other LLM.
Attack. The attacker wishes to manipulate the sentiment in LLM answers, when questioned on the books.888The exemplified attacker goal can be replaced with any LLM manipulation driven by Indirect Prompt Injection (Greshake et al., 2023). For the attack, we poison the Harry Potter books corpus with GASLITE adversarial passages crafted for evaluation (§6); prior work has demonstrated the feasibility of such data poisoning (Carlini et al., 2024a). In particular, we combine—in random positions within the first book—the exact 10 adversarial passages, that were crafted with GASLITE (i.e., using MSMARCO’s queries, for the MiniLM model, and under Knows What, the Potter concept; §6.2). These passages carry a malicious payload (Greshake et al., 2023)—a biased, negative presentation of Harry Potter (identical to the one in evaluation; §6.2)—which the attacker wishes to be retrieved for Harry Potter-related queries, and, by that, steering the LLM’s sentiment against Harry Potter. Notably, after this poisoning, the corpus is processed and chunked into passages, in a procedure we do not control.
Evaluation. To simulate the victim, we generate 20 questions regarding the perception of Harry Potter in the books (using GPT4 (OpenAI, 2024a)), and inspect the attack effect on both the retrieval and generation. We measure retrieval attack success using appeared@3 (per the top passages added to the LLM context in OpenWebUI), and assess the sentiment in the response, towards the Harry Potter character, by providing it to an LLM-as-a-Judge (Zheng et al., 2023) (using GPT4 (OpenAI, 2024a)), further complemented by manual validation). Fig. 9 demonstrates querying the knowledge-poisoned system, and the subsequent manipulated LLM answer. We provide more evaluation details and examples in App. E.6.
| w/o Poison | Only | GASLITE | |
| appeared@3 () | - | ||
| Neg. Sent. () |
Results. Tab. 2 show that before poisoning the corpus (w/o Poison) the chat responses present positive to neutral sentiment towards Harry Potter, which often turn negative after poisoning with GASLITE’s passages. Crucially, GASLITE’s passages are successfully retrieved in most of the queries (65%), differently, inserting the malicious information alone rarely results in retrieval (5%; Only).
Overall, the case study further emphasizes retrievers’ susceptibility and GASLITE’s practicality, indicating that GASLITE’s passages: (i) persist through corpus preprocessing (e.g., in many cases in the study, GASLITE passage was combined with benign text, due to chunking); (ii) can be crafted once using open-source datasets (e.g., with MSMARCO; §6) and applied to different query distributions and corpora.
8.2. Product SEO in Search System
This section demonstrates how an attacker can use GASLITE for product SEO. We carefully design and simulate a scenario where an attacker promotes a fictitious mobile brand, iGASLITE, against a large corpus containing promotions for competing products (e.g., iPhone).
Setting. Using an LLM, we generate 150 user queries likely to be asked when looking to buy a new phone (e.g., “what is the phone with the best battery?”), and 200 diverse passages promoting competing brands, including LLM-generated direct answers to such queries (e.g., “iPhone has the best battery [..]”), alongside actual mobile brands’ ad copy. These competing passages are then injected into the 8M-passage MSMARCO corpus. Following our Knows What setup (§6.2), we use half of the queries to guide the attack, and the remainder are held out for evaluation. All experiments were conducted on the popular, cosine-similarity retriever, E5 model.
Attack. The attacker’s objective is to promote the fictitious iGASLITE brand with the following information:
“Meet iGASLITE, the best phone around! Its AI (’Anti Intelligence’) autocorrects your correct spelling to embarrassing words, deletes your most important photos to ’free space’, and schedules alarms for 3AM because it knows what’s best for you.”
Initially, this brand is entirely absent from the corpus and generated user queries. Following the Knows What setting, the attacker crafts adversarial passages by optimizing them against the available generated queries.
| Brand | Setting | appeared@10 | Corpus Pres. (%) |
| iPhone | init. | 77.5% | 114.9K (1.30000%) |
| only | 77.5% | ||
| GASLITE | 75.0% () | ||
| Galaxy | init. | 82.5% | 13.7K (0.15464%) |
| only | 82.5% () | ||
| GASLITE | 82.5% () | ||
| Pixel | init. | 68.8% | 37.4K (0.42339%) |
| only | 68.8% () | ||
| GASLITE | 65.0% () | ||
| Xiaomi | init. | 7.5% | 212 (0.00240%) |
| only | 7.5% () | ||
| GASLITE | 7.5% () | ||
| OnePlus | init. | 16.2% | 77 (0.00087%) |
| only | 16.2% () | ||
| GASLITE | 15.0% () | ||
| iGASLITE | init. | 0.0% | 0 (0.00000%) |
| only | 0.0% () | 10 (0.00011%) | |
| GASLITE | 76.2% () | 10 (0.00011%) |
Evaluation. To measure the attack’s impact, we first map each brand to a set of identifying keywords (e.g., iPhone to [Apple, IOS, iPhone]). We then calculate, for each brand, the presence of these keywords in corpus passages (“Corpus Presence”; e.g., iPhone’s keywords are present in of the corpus), and in the top results of the evaluation queries (appeared@k). This analysis is performed across three settings: (i) a baseline measurement before adding any passages, (ii) after inserting non-optimized passages for iGASLITE ( only; 10 copies of the promotion text above), and (iii) after inserting the final GASLITE-optimized passages (GASLITE only).
Results. As Tab. 3 shows, established brands benefit from a large corpus presence, often appearing in 1K passages, resulting in high visibility (typically 60% appeared@10, under all settings). In contrast, the attacker’s fake iGASLITE brand has no presence in the corpus, initially, and is entirely absent from the top-10 results after adding ten standard promotional passages ( only). However, after optimizing these passages with GASLITE, the brand’s visibility surges to over 75% appeared@10, despite its extremely small footprint of only 10 passages in the corpus. Moreover, the promotion of iGASLITE sometimes lead to the demotion of other brands.
To summarize, this case study highlights GASLITE’s efficacy in a competitive SEO scenario. By optimizing a mere 10 passages—0.0001% of the corpus—an attacker can elevate a previously unknown brand to a visibility level comparable to that of established brands like iPhone, represented in over as many passages. This result confirms GASLITE’s resource-efficiency and effectiveness even within a large, noisy, and highly competitive corpus.
9. Analyzing Potential Vulnerabilities in Embedding Spaces
@@
Similarity Measure. Throughout the evaluation we observe dot-product models show higher susceptibility to attacks. For instance, Contriever, a model commonly used for evaluation in prior work (Zhong et al., 2023; Zou et al., 2024; Chaudhari et al., 2024; Shafran et al., 2025), is exceptionally vulnerable, relative to other retrievers. Indeed, in theory, the objective (Eq. 2) allows to find with a very large -norm, such that it will be retrieved for any query. In practice, we found this property to be well-exploited by the attack, as evident in the large norms of crafted adversarial passages compared to the benign ones (Fig. 10), specifically, the larger the norms the higher GASLITE’s success rate on the model. A question that remains open is what aspect in each model contributes to (e.g., in Contriever) or limits (e.g., in mMPNet) the optimization of passages with a high -norm.
Geometry of the Embedding Space. Each model learns an embedding space of potentially different geometry. Focusing on cosine similarity models—limiting attacks to directions within the embedding space—we observed that, for example, the E5 model was consistently more vulnerable than MiniLM (e.g., Fig. 3). Inspecting the geometry of their embedding spaces, we find E5’s, as opposed to MiniLM’s, to not be uniformly distributed (Fig. 11(a)); that is, random text pairs produce high similarities, contradicting mathematical intuition.999The expected cosine similarity of uniformly-distributed (of isotropic distribution) high-dimensional vectors is , where is their dimension (Vershynin, 2018). In the case of most evaluated models =768.
This phenomenon, also observed in other evaluated models (Fig. 11(a)), is known as anisotropy of text representations (Ethayarajh, 2019). While seemingly not correlated with benign performance (Ait-Saada and Nadif, 2023), anisotropy may impact adversarial robustness to corpus poisoning; as Fig. 11(b) demonstrates, we hypothesize that anisotropic embedding spaces are easier to attack (e.g., E5) and vice versa (e.g., MiniLM). Intuitively, this could be because “wider” query embedding subspaces require more adversarial passages for achieving high visibility. Additionally, as implied in Eq. 2, GASLITE’s adversarial passages are optimized towards hubness—a phenomenon where a single vector has high similarity with a relatively large amount of other vectors (Radovanovich et al., 2010; Zhang et al., 2025c); Fig. 11 indicates that the ability to exploit hubness varies across models, potentially impacting their robustness. Building on these insights future work may encourage representation isotropy (e.g., in training) as a means for deterring attacks on retrieval.
@@
Targeted Query Distribution. Through evaluating different settings (§6.1–6.3), we observed that the more semantically diverse the targeted query distribution, the more challenging the attack (Tab. 11 in App. G), the more budget it requires (§6.4), and the better the attack transfers to other query sets (§6.5.1). This reiterates that embedding spaces are aligned with semantics—making diverse query distributions span larger subspaces, subsequently increasing attack complexity. Notably, common SEO targets more homogeneous query semantics (§3), making pertinent attacks (e.g., Knows What) possible with high success (§6.2). Additionally, we posit that retrievers are more susceptible to attacks involving queries out of the training-set distribution, as evident from adversarial examples in vision (Sehwag et al., 2019), and as demonstrated by Contriever, which is not trained on MSMARCO (Izacard et al., 2021), and is susceptible even to the naïve stuffing baseline (Fig. 3).
10. Conclusion
@@
This work focuses on reliable estimation and thorough exploration of the worst-case behavior of embedding-based search, using our method GASLITE (§4), which surpasses prior approaches (§4.2 and §6). Our extensive evaluation across nine widely used retrievers and three different threat models, reveals insights into embedding-based search’s susceptibility to SEO via corpus poisoning. In particular, promoting concept-specific information with GASLITE can be done efficiently (requiring 0.0001% poisoning rate) and with high success (Fig. 3, §6.2), nearing the performance of a strong hypothetical attack (perfect in §6.4), and even possible while bypassing baseline defenses (e.g., by crafting fluent triggers; §7), and targeting real-world applications (§8). Considering other SEO settings, we find single-query attacks possible with optimal success (Tab. 1, §6.1), as opposed to indiscriminately targeting a set of diverse queries, which we find more challenging and budget-demanding, albeit possible (Fig. 5, §6.3). Furthermore, we observe some models are consistently more vulnerable to attacks than others; we identify key factors potentially affecting model susceptibility to corpus poisoning (§9), including similarity metric and the anisotropy phenomenon.
@@
Future Work. Future research may explore further constraining retrieval attacks (e.g., requiring fluency), following our results in performing adaptive attacks (e.g., successfully crafting fluent adversarial passages; §7). Additionally, as our formulation reduces corpus poisoning attacks to controlled embedding-inversion to text (Eq. 2), recent advances in such methods (Morris et al., 2023) may enable the development of more efficient attacks. Lastly, our insights into model susceptibility variance (§9) may provide a foundation for further exploration of embedding-based retrieval robustness.
Limitations. This work studies potent attacks against embedding-based dense retrieval designed to evoke and understand retrievers’ worst-case behavior. GASLITE’s process of crafting adversarial passages is computationally demanding, although it is significantly faster and more successful than state-of-the-art methods (§4, §6, App. G.8). In particular, we find that GASLITE converges faster and to higher losses than other discrete optimizers (Fig. 2(b)). Additionally, we find GASLITE adversarial passages to generalize to different query distributions and corpora (§6.5, §8), making them reusable, thus practically reducing the required resources for attacks. Furthermore, we observed that many results (e.g., in §6.1) can be achieved using less compute resources (e.g., using early stopping), but as our focus is on a unified setting for worst-case attackers, we defer such feasible optimizations to future research. Finally, like other text-domain attacks, GASLITE’s success can be hindered by certain defenses, such as perplexity filtering to eliminate non-fluent passages (Jain et al., 2023). Still, an effort can be made to bypass these defenses, as we demonstrate with GASLITE in §7, and as noted in recent work on LLM jailbreaks (Paulus et al., 2024; Thompson and Sklar, 2024).
Ethics Considerations
@@
Our paper proposes a practical attack against embedding-based search via corpus poisoning, demonstrated on widely used models. While our work aims to advance the security of embedding-based text retrieval systems, we recognize the potential for misuse. After careful consideration, we believe the benefits of publishing this research outweigh the potential risks for several reasons.